# Does VLM Classification Benefit from LLM Description Semantics?

### v1

@@ 5 Conclusion
In our study, we found that VLM Classification performance does indeed benefit from LLM description semantics - if the descriptions are correctly selected. For that, we present a new, training-free approach that assigns semantically meaningful descriptions based on feedback from the VLM embedding space. We show that descriptions assigned in this way do have discriminative power on their own, i.e. when evaluated in our proposed setup that has no classname ensembling. These description assignments improve the performance in image classification tasks - with and without classname ensembling. Furthermore, this evaluation setup helps to differentiate performance gains from genuine semantics versus gains from ensemble effects. Hopefully, our findings inspire future work on VLMs and model interpretability.
@@ A.1 Limitations and Ethical Considerations
Currently, most of the methods in this domain, including ours, work with a fixed pool of descriptions $\mathcal{P}$ from LLMs. Although our method finds precise and meaningful descriptions for better performance, it would be interesting to give such selection feedback to LLMs and let them refine the $\mathcal{P}$ in an agentic fashion. Furthermore, due to resource limitations, we could not test our method on a more diverse and scalable dataset. However, we anticipate it will still perform well. We are not aware of any ethical considerations, currently.
@@ A.2 Different choices of size $k$ and $n$
Table 5 ablates the number $k$ of preliminary labels per test image and $n$, the number of images per class available in the selection set for S. Results were obtained in the classname-included evaluation setup on the CUB200 Dataset with $5$ selected descriptions from the DClip description pool. Increasing the number of selection samples in $S$ leads to increasing performances. For the CUB dataset, the optimal $k$ value is empirically found to be $k=3$ as it has a right trade-off between a diverse set of candidates that is still small enough to find distinctive descriptions inside. This shows that even in a few-shot regime with only $5$ selected descriptions per candidate class baselines such as  (Esfandiarpoor and Bach 2024) and  (Menon and Vondrick 2022) can be exceeded (cf.  Table 3 in the main text).
@@
Table 5: Hyper-params comparison on CUB200 dataset for ViT-B/32 backbone $k$ defines the number of preliminary labels per test image to consider, and $n$ denotes the number of images per class available in the selection set for S.
A.3 Choice of heuristics.
Table 6 shows the results when different $\Phi(\cdot,\cdot)$ is applied. On datasets that work well under classname-free scenario (see Figure 3), i.e. shows the characteristic bump for $w_{cls}$ in $[0,10]$, the maxmean heuristic works better. On datasets where this bump is absent (ImageNet, Food101, and Oxford Pets), the maxmin heuristic performs better. This can come from the fact that finding descriptions that distinguish one preliminary class from all others simultaneously is challenging for these datasets. Since the maxmin heuristic focuses on maximizing the smallest similarity difference, future works should explore pairwise comparisons to mine descriptions for these datasets.
Heuristics $\Phi(\cdot,\cdot)$ ImageNetV2 ImageNet CUB200 EuroSAT Places365
maxmean 60.13 60.09 55.16 49.30 41.60
maxmin 61.05 60.60 51.95 49.58 40.26
Table 6: Choice of the selection heuristics $\Phi(\cdot,\cdot)$
A.4 Ablation for description assignments with $\mathcal{P}$ from DCLIP.
@@
LLM assignments, global eval 13 61.99 55.09 51.79 43.31 39.91 43.09 62.97
LLM assignments, local top-k eval 13 61.99 55.06 51.83 43.29 39.87 43.09 62.86
random assignments, global eval 13 61.88 54.86 52.00 41.46 38.98 43.24 62.81
random assignments, local top-k eval 13 61.91 54.86 51.97 41.41 39.02 43.30 62.82
@@
A.5 Results for ViT-L/14 CLIP Backbone
We also show evaluation results on the ViT-L/14 CLIP backbone, as it is commonly used for an additional comparison. One can see a consistent pattern as we showed in the main paper for the ViT-B/32. Cf.  Table 9 and  Table 8.
@@ A.5 Results for ViT-L/14 CLIP Backbone
Table 8: Image classification with classname-included descriptions for ViT-L-14 backbone.
Source of $\mathcal{P}$ Description Assignment Max #desc. $\downarrow$ ImageNet ImageNetV2 CUB200 EuroSAT Places365 Food101 Oxford Pets DTD Flowers102
DCLIP LLM assignments 12 73.66 68.05 63.07 56.85 41.54 90.80 93.32 53.56 74.84
@@
Contrastive LLM assignments 40 73.57 68.04 62.82 55.91 41.85 90.82 93.73 54.84 75.09
Contrastive random assignments 40 73.07 67.96 63.38 56.39 40.62 90.78 93.24 52.87 74.86
@@
Table 9: Image classification in classname-free setup for ViT-L-14 backbone. Maximum value for the best performing $w_{cls}$ per cell.
A.6 Ablating ambiguous context size $k$
Table 10 ablates parameter $k$, which denotes the number of ambiguous classes considered per test image $x_{i}$. Different datasets behave differently under a changed $k$. Interestingly, datasets with challenging description behavior - namely Food101, ImageNet and ImageNetV2- do not show a high sensitivity to a varying $k$. The best and worst performances for these datasets do not differ much, although a tendency for slightly better performances for higher values of $k$ can be observed. The remaining datasets with favorable description behavior are much more sensitive to the chosen values of $k$: CUB200 and EuroSAT [^1] In this evaluation $n=2000$ vs. $n=1000$ in Table 3. The reduced bias explains the stronger results. perform best for low values of $k$ while DTD, Flowers102 and Places365 perform best for medium to high values of $k$. However, for any chosen value of $k$, LLM assignments are surpassed significantly for these datasets. Similar results are confirmed in Table 11 for the ViT-L/14 backbone where Food101, ImageNet, and ImageNetV2 classification performance reacts insensitively to $k$. Interestingly, for ViT-L/14, high $k$ values benefit the classification performance, whereas for ViT-B/32, low $k$ values yielded the best results.
Besides that, it is remarkable that the language-maxing accuracy denoted in parentheses closely matches or even surpasses [^2] In case of the ViT-L/14 backbone. the ensembling accuracy, although it cannot make use of the smoothing ensembling effect that can work with random descriptions (see Section A.7). This points to the distinctive quality of the selected descriptions.
@@ A.6 Ablating ambiguous context size $k$
A.7 Evaluating Beyond Ensembling
@@
Classification Accuracies for ViT-B/32
Description Pool Assignment Evaluation Mode ImageNet ImageNetV2 CUB200 EuroSAT Places365 DTD Flowers102
@@
Table 12: Comparison of semantic-sensitive maxing evaluation for both assignment types within the DCLIP description pool. Evaluation happens in the classname-containing, conventional scenario to compare the classname ensembling effect to the non-ensembling maxing evaluation. Ensembling of LLM-assigned descriptions is referenced as a baseline. Colored differences between maxing evaluation and baseline performance are displayed. For LLM-assigned descriptions, maxing performance falls short of the ensembling baseline performance in $6/7$ cases and exceeds it in no case. Contrary to this, the maxing evaluation for the selected assignments exceeds ensembling baseline performance in $5/7$ cases with considerable gains. Thus, the enhanced semantic distinctiveness of the selected descriptions enables to surpass the classname ensembling effect.
A.8 Construction of the description pool $\mathcal{P}$ from existing method.
@@
A.9 Comparison to prior works in the classname-free setting
@@
A.10 Chosen selection sizes $n$
@@
A.11 Examples of selected descriptions
@@
A.12 Example of a description pool P

### v2

@@ 5 Conclusion
This study found that VLM Classification performance does indeed benefit from LLM description semantics - if the descriptions are correctly selected. For that, we present a new, training-free approach that assigns semantically meaningful descriptions based on feedback from the VLM embedding space. We show that descriptions assigned in this way have discriminative power on their own, i.e. when evaluated in our proposed setup without classname ensembling. These description assignments improve the performance in image classification tasks - with and without classname ensembling. Furthermore, this evaluation setup helps to differentiate performance gains from genuine semantics from gains from ensemble effects. Hopefully, our findings inspire future work on VLMs and model explainability.
Acknowledgements
This project has been supported by the German Federal Ministry for Economic Affairs and Climate Action within the project “NXT GEN AI METHODS – Generative Methoden für Perzeption, Prädiktion und Planung”, the German Research Foundation (DFG) project 421703927, Bayer AG, and the bidt project KLIMA-MEMES. The authors gratefully acknowledge the Gauss Center for Supercomputing for providing compute through the NIC on JUWELS at JSC and the HPC resources supplied by the Erlangen National High Performance Computing Center (NHR@FAU funded by DFG).
@@ A.1 Limitations and Ethical Considerations
Currently, most of the methods in this domain, including ours, work with a fixed pool of descriptions $\mathcal{P}$ from LLMs. Although our method finds precise and meaningful descriptions for better performance, it would be interesting to give such selection feedback to LLMs and let them refine the $\mathcal{P}$ in an agentic fashion. We are not currently aware of any ethical considerations.
@@ A.2 Different choices of size $k$ and $n$
Table 4 ablates the number $k$ of preliminary labels per test image and $n$, the number of images per class available in the selection set for S. Results were obtained in the classname-included evaluation setup on the CUB200 Dataset with $5$ selected descriptions from the DClip description pool. Increasing the number of selection samples in $S$ leads to increasing performances. For the CUB dataset, the optimal $k$ value is empirically found to be $k=3$ as it has a right trade-off between a diverse set of candidates that is still small enough to find distinctive descriptions. This shows that even in a few-shot regime with only $5$ selected descriptions per candidate class baselines such as  (Esfandiarpoor and Bach 2024) and  (Menon and Vondrick 2022) can be exceeded (cf.  Table 3 in the main text).
@@
Table 4: Hyper-params comparison on CUB200 dataset for ViT-B/32 backbone. $k$ defines the number of preliminary labels per test image to consider, and $n$ denotes the number of images per class available in the selection set for S.
A.3 Ablation for description assignments with $\mathcal{P}$ from DCLIP.
@@
LLM 13 61.99 55.09 51.79 43.31 39.91 43.09 62.97
LLM (local-k) 13 61.99 55.06 51.83 43.29 39.87 43.09 62.86
random 13 61.88 54.86 52.00 41.46 38.98 43.24 62.81
random (local-k) 13 61.91 54.86 51.97 41.41 39.02 43.30 62.82
@@
A.4 Results for ViT-L/14 CLIP Backbone
We also show evaluation results on the ViT-L/14 CLIP backbone, as it is commonly used for an additional comparison. One can see a consistent pattern as we showed in the main paper for the ViT-B/32. See Table 7 and Table 6.
@@ A.4 Results for ViT-L/14 CLIP Backbone
Table 6: Image classification with classname-containing descriptions for the ViT-L-14 backbone. The same parameters as in Table 3 were used.
Source of $\mathcal{P}$ Assignment Max #desc. $\downarrow$ ImageNet ImageNetV2 CUB200 EuroSAT Places365 Food101 Oxford Pets DTD Flowers102
DCLIP LLM 12 73.66 68.05 63.07 56.85 41.54 90.80 93.32 53.56 74.84
@@
Contrastive LLM 40 73.57 68.04 62.82 55.91 41.85 90.82 93.73 54.84 75.09
Contrastive random 40 73.07 67.96 63.38 56.39 40.62 90.78 93.24 52.87 74.86
@@
Table 7: Image classification in classname-free setup for the ViT-L-14 backbone. Maximum value for the best performing $w_{cls}$ per cell. The same parameters as in Table 1 were used.
A.5 Ablating $k$
Table 8 ablates parameter $k$, which denotes the number of ambiguous classes considered per test image $x_{i}$. Different datasets behave differently under a changed $k$. Interestingly, datasets with challenging description behavior - namely Food101, ImageNet and ImageNetV2 - do not show a high sensitivity to a varying $k$. The best and worst performances for these datasets do not differ much, although a tendency for slightly better performances for higher values of $k$ can be observed. The remaining datasets with favorable description behavior are much more sensitive to the chosen values of $k$: CUB200 and EuroSAT perform best for low values of $k$ while DTD, Flowers102, and Places365 perform best for medium to high values of $k$. However, for any chosen value of $k$, LLM assignments are surpassed significantly for these datasets. Similar results are confirmed in Table 9 for the ViT-L/14 backbone where Food101, ImageNet, and ImageNetV2 classification performance reacts insensitively to $k$. Interestingly, for ViT-L/14, high $k$ values benefit the classification performance, whereas for ViT-B/32, low $k$ values yielded the best results.
Besides that, it is remarkable that the language-maxing accuracy denoted in parentheses closely matches or even surpasses [^1] In case of the ViT-L/14 backbone. the ensembling accuracy, although it cannot make use of the smoothing ensembling effect that can work with random descriptions (see Section A.6 for more information). This points to the distinctive quality of the selected descriptions.
@@ A.5 Ablating $k$
A.6 Evaluating Beyond Ensembling
@@
Pool Assignment Evaluation ImageNet ImageNetV2 CUB200 EuroSAT Places365 DTD Flowers102
@@
Table 10: Comparison of semantic-sensitive maxing evaluation for both assignment types within the DCLIP description pool for ViT-B/32. Evaluation happens in the classname-containing, conventional scenario to compare the classname ensembling effect to the non-ensembling maxing evaluation. Ensembling of LLM-assigned descriptions is referenced as a baseline. Colored differences between maxing evaluation and baseline performance are displayed. For LLM-assigned descriptions, maxing performance falls short of the ensembling baseline performance in $6/7$ cases and exceeds it in no case. Contrary to this, the maxing evaluation for the selected assignments exceeds ensembling baseline performance in $5/7$ cases with considerable gains. Thus, the enhanced semantic distinctiveness of the selected descriptions enables to surpass the classname ensembling effect.
A.7 Construction of the description pool $\mathcal{P}$ from existing method.
@@
A.8 Comparison to prior works in the classname-free setting
@@
A.9 Chosen selection sizes $n$
Per dataset, the sample sizes $n$ that are displayed in Table 11 were utilized for the experiments of Table 1 and  Table 3.
@@ A.9 Chosen selection sizes $n$
A.10 Varying the LLM.
LLMs ImageNet ImageNetV2 CUB200 EuroSAT Places365 DTD Flowers102
Contrastive Llama3-70B 61.68 54.64 53.49 36.33 41.33 79.87 64.45
w/ Our Selection 63.30 55.24 56.27 58.57 43.65 48.09 68.61
Contrastive GPT3.5 61.65 54.69 53.83 35.38 40.79 44.10 66.29
w/ Our Selection 63.15 55.09 56.21 44.43 43.64 80.78 68.48
A.11 Varying the VLM.
Table 13 demonstrates that our algorithm functions effectively within the ALIGN VLM embedding space (Jia et al. 2021). This further illustrates the versatility of our method.
ALIGN ImageNet ImageNetV2 CUB200 EuroSAT Places365 DTD Flowers102
ALIGN 59.83 54.59 37.54 29.62 39.89 55.00 54.26
+LLM assigned 59.87 54.84 37.25 29.37 40.24 57.29 55.28
+randomly assigned 59.83 54.57 37.14 29.14 39.89 57.18 55.46
w/ Our Selection 60.26 55.29 40.32 31.74 42.47 59.79 55.91
Table 13: Performance in the classname-free evaluation setup using the ALIGN VLM.
A.12 Examples of selected descriptions
@@
A.13 Example of a description pool P
@@
A.14 Distribution of Distinctiveness Scores
Figure 4 shows the ranked distribution of distinctiveness scores obtained by the training-free method in $\emph{S}^{+}_{i}$ of Algorithm 1. For one randomly chosen image per dataset and one randomly chosen ambiguous class, all positive values of $\overline{\mathrm{diff}}_{a,a\prime\in\mathcal{A}}^{d}$ are displayed, sorted by their rank. These values correspond to the distinctiveness of descriptions in $S^{+}$. Notably, the ranked distribution of the distinctiveness scores appears to follow a Pareto distribution: A few top-ranked descriptions score substantially higher than the rest, while most descriptions score significantly lower and are closer to each other, which is characteristic of a Pareto distribution. These highest-scoring descriptions offer the highest distinctiveness according to the available samples in $S$ w.r.t. $a$. Since only relatively few descriptions yield high distinctiveness scores, this concise set of highly distinctive descriptions can be captured by a low value of $m$. Thus, selecting the top-$m$ scoring descriptions via $m=5$, already captures large parts of the steeply declining highly distinctive descriptions. This explains why already $5$ selected descriptions in the DCLIP description pool bring substantial performance gains, as seen in Table 3 and Table 6.
Figure 4: Distinctiveness scores of randomly chosen images obtained by the training-free approach presented in Section 3.2. Distinctiveness scores $\overline{\mathrm{diff}}_{a,a\prime\in\mathcal{A}}^{d}=\frac{1}{k-1}\sum_{a\prime\in\mathcal{A}}\mathrm{diff}_{a,a^{\prime}}^{d}=\bar{s}_{a,d}-\bar{s}_{a^{\prime},d}$ where $\mathrm{diff}_{a,a^{\prime}}^{d}=\bar{s}_{a,d}-\bar{s}_{a^{\prime},d}\geq 0$. Used Parameters: $k=3$, $m=5$, $n=\text{maximal}$, pool = DCLIP. See Section A.14 for a concise discussion.